OpenStudy (ajprincess):

Pls help:) Show that \(x^3+3xy^2=1\) gives a solution of the differential equation \[2xy\frac{dy}{dx}+x^2+y^2=0\] on the interval 0

OpenStudy (amistre64):

y^2 = (1-x^3)/(3x)

OpenStudy (amistre64):

but thats prolly inconsequential

OpenStudy (ajprincess):

I have showd that it is a solution.but i dnt knw how to show the interval part.

OpenStudy (amistre64):

D[x^3 + 3x y^2 = 1] D[x^3] + D[3x y^2] = D[1] 3x^2 x' + D[3x y^2] = 0 3x^2 x' + (3x D[y^2] + D[3x] y^2) = 0 3x^2 x' + (3x 2y y' + 3x' y^2) = 0 assumeing its wrt.x; x'=1 3x^2 + (3x 2y y' + 3y^2) = 0 solve for y'

OpenStudy (zarkon):

you don't need to solve for y'

OpenStudy (amistre64):

2xy not equal to 0

OpenStudy (zarkon):

just factor out a 3

OpenStudy (amistre64):

but i like solving for y' its my favorite part :)

OpenStudy (ajprincess):

:) but how do i show that it lies in that interval?

OpenStudy (amistre64):

if 2xy not equal to 0, then the rest of it looks fine to me recall that we would need to divide off that 2xy to get a proper solution for some reason

OpenStudy (amistre64):

but i spose that involves finding a y=_____

OpenStudy (amistre64):

y^2 = (1-x^3)/(3x) y =+- sqrt( (1-x^3)/(3x)) so as long as 1-x^3 not equal zero .....

OpenStudy (amistre64):

since x = 1 is bad, and x=1 is not in the interval ... its good

OpenStudy (ajprincess):

i am really sorry. i am new to differential equation. so this interval part is still confusing me.:(

OpenStudy (amistre64):

\[ky'+ay=b\] has a solution (unique solution) if k not equal to 0, why?

OpenStudy (amistre64):

\[y'=\frac{b-ay}{k}\]

OpenStudy (ajprincess):

if k is equal to zero then it becomes undefined which leaves with no solution. is that the reason @amistre

OpenStudy (amistre64):

thats a rather simplistic reason yes :) i cant say that im confident for any other reasons ive come across

OpenStudy (amistre64):

a solution has to be continuous along an interval, and if the interval has an undefinable point ... then continuity is in question

OpenStudy (ajprincess):

oh. so i need to find two points where it continuity is in question. in this case when x=0 y becomes undefined and when x=1 it becomes 0. simply if we get infinity or 0 thn that is the point where continuity ends. am I right?

OpenStudy (amistre64):

you would need to find at most, 1 point, where continuity is in question we need to look at all the possible situations where we get a "bad math" issue. x^3+ 3x y^2 = 1 , in the original equation there is no problem \[2xy\frac{dy}{dx}+x^2+y^2=0\] \[\frac{dy}{dx}+\frac{x^2+y^2}{2xy}=0\]we have an issue when 2xy = 0; so when x=0, or y=0 is problematic we know that x=0 is not in the interval, so lets determine of y=0 happens ... x^3+ 3x y^2 = 1 , solve for y, or y^2 even \[y^2=\frac{1-x^3}{3x}\] since there is no x=0 in the interval, that leaves us with 1-x^3 = 0 to eliminate

OpenStudy (amistre64):

we have 2 solutions for y, which is what i initially drew up \[y=\sqrt{\frac{c-x^3}{3x}}\] \[y=-\sqrt{\frac{c-x^3}{3x}}\] at x=c we therefore have 2 equally valid solutions and cannot determine a unique solution

OpenStudy (amistre64):

since we know the particular solution is c=1 .....

OpenStudy (amistre64):

so, in the interval: 0< x < 1 we CAN define a unique solution for all values of x with the given solution, is what i get from it all

OpenStudy (ajprincess):

I get it now. Thanxxx a lot for ur help.:D

OpenStudy (amistre64):

youre welcome :)

OpenStudy (ajprincess):

@amistre sorry to trouble u. if i only show that x^3+3xy^2=1 gives a solution the d.e is it enough? or should I show what happens when x<0 or x>1?

OpenStudy (ajprincess):

@amistre64

OpenStudy (ajprincess):

@Zarkon

OpenStudy (amistre64):

since they give you the interval: 0<x<1 it is always good practice to test if the interval is valid. The textbooks tend to mention that after they teach this portion of the material that it is assumed that the solutions exist within their proper intervals

OpenStudy (amistre64):

i believe one good way to consider this interval of existence debacle is to relate it to the interval of convergence of a power series.